OpenStudy (anonymous):

The hypotenuse of a right triangle is 7 inches, and one of the legs is √13 inches. Find the length of the other leg.

OpenStudy (anonymous):

\[7^2=x^2+\sqrt{13}^2\] \[49=x^2+13\] \[36=x^2\] \[x=6\]

OpenStudy (anonymous):

first line is pythagoras \[h^2=a^2+b^2\] the plug in the numbers and solve for the one you do not know

OpenStudy (anonymous):

A trig solution:\[7 \text{Cos}\left[\text{ArcSin}\left[\frac{\sqrt{13}}{7}\right]\right]=6 \]